the probability that the 13th day of a randomly chosen month is a friday?

Grade:

4 Answers

Chetan Mandayam Nayakar
13 years ago

any randomly chosen day has the same probability, namely 1/7 of being any day of the week from monday to sunday. therefore the answer is 1/7.

'Win exciting gifts by answering the questions on Discussion Forum.

Uddesh Nath
7 years ago
Probability of choosing any month out of 12 month =1/12 Probability that the 13th day is Friday if the 1st day say starts with Sunday =1/7 Thus required probability =1/12×1/7 =1/84
srikanth goud
6 years ago
probability of selecting month = 1/12 and probability of selecting friday = 1/7.so,required probability = 1/12 x 1/7 = 1/84
You know who
5 years ago
There are 12 months and 7days so permutations can be 12*7=84 . Therefore the probability of getting a Friday will be 1/84